TANGENTS AND NORMALS
at what points of the curve y = x^2 does the tangent make an angle of 45 degree with the x axis.

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

Slope = m
m = tan45 = 1
y = x^2
y' = 2x
2x = 1
x = \frac{1}{2}
Put in curve:
y = \frac{1}{4}